9. Given the following information, find fixed costs: a. Total sales, $104,672 profit, $18.000; variable rate, 42
Question 9 10. Thanks for your help
9. Given the following information, find fixed costs: a. Total sales, $104,672 profit, $18.000; variable rate, 42
h Profit, $12.000: number of customers, 32392; variable cost per unit. c. Sales price per unit, $1460; profit, $34,000, number of customers. price per unit, $18.40: number of custom- $4.63; sales price per unit, $10.34 26,712; variable rate, 35 ers, 26,549: profit, $33,000 d. Contribution rate, 65; sales 10. Given the following information, find profit xed costs, $82,449.40; total sales, $167,543.20: variable cost $55,629.60 $193,764.40; sales price per unit, $17.60 $75,919.70; sales price per unit, $16.22 b. Variable rate, 4 number of customers, 26.412; fixed costs, c. Total sales, $190,830.66; variable cost per unit, $5.64: fixed costs,
Answer #1

9)

A) Total sales = variable cost + fixed cost + profit

Form the above

104,672= 104,672× 0.42 + fixed cost +18,000

Fixed cost = $ 42,709.76

=~ $ 42,710

B)

Total sales = variable cost + fixed cost + profit

Form the above

32,392 × 10.34 = 32,392× 4.63 + fixed cost + 12,000

Fixed cost = $172,958.32

C) Here given variable cost in percentage on sales

Total sales = variable cost + fixed cost + profit

Form the above

26,712×14.60 = 26,712 × 14.60× 0.35 + fixed cost +34,000

Fixed cost = 219,496.88

D)here given contribution it represents that the portion of sales revenue that is not consumed by variable cost and contribute to coverage of fixed cost.

Contribution = sales price × contribution ratio

= 18.40 ×0.65 =11.96

Formula Profit = contribution - fixed cost

18000 =26,549× 11.96 - fixed cost

Fixed cost =299,526.04

10)

A)Profit = Total sales - variable cost - fixed cost

= 167,543.20 -55,629.60-82,449.40

Profit = 29,464.2

B) Profit = total sales - total variable cost - fixed cost

= 26,412 ×17.60 - 26,412 × 17.60× 0.4 - 193,764.40

Profit = 85,146.32

C) Here we calculate as per given below

Sales(units) = sales in amounts/sale price per unit

= 190,830.66/16.22 =11,766 units

Contribution per unit = sale price per unit - variable cost per unit

= 16.23-5.64 = 10.59

Profit =total contribution - fixed cost

= 11,766× 10.59 - 75,919.70 = 48,682.24